OpenStudy (anonymous):

Auditions for a play in New York were held with 100 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? 44 36 76 60

OpenStudy (anonymous):

im msellier sister need help for this please

OpenStudy (anonymous):

anyone please?

OpenStudy (mathmate):

Let M=number of men 100-M = number of women Set up equation from " If twice the number of men less three times the number of women is equal to 20" 2M-3(100-M)=20 Now use the equation to solve for M and subsequently find the number of men and women.

OpenStudy (mathmate):

Make sure the result is an integer (i.e. whole number)!

OpenStudy (anonymous):

64 ?

OpenStudy (mathmate):

For number of men?

OpenStudy (anonymous):

2M-3(100-M)=20 2M - 300 + 3M =20 5M-300= 20 add 300 each side 5M = 320 M = 64

OpenStudy (anonymous):

? @mathmate ?

OpenStudy (mathmate):

The answer is good!

OpenStudy (mathmate):

Do check what the question is asking!

OpenStudy (anonymous):

oh so then substitute 64 for M ?

OpenStudy (mathmate):

In retrospect, I should have set up: W=number of women, and (100-W)=number of men... But either way, you can get the right answer.

OpenStudy (mathmate):

The question asks for the number of women who auditioned.

OpenStudy (anonymous):

im lost?

OpenStudy (mathmate):

If there are 100 persons who auditioned, out of which 64 are men. How many women auditioned?

OpenStudy (anonymous):

oh lol i see yeah so 36 women

OpenStudy (anonymous):

?

OpenStudy (mathmate):

Yes, that is correct.

OpenStudy (anonymous):

thank you alot msellier sister :)

OpenStudy (mathmate):

yw!